Out of a sample of 325: 274 said the were in favor of campaign finance reform, 98 were in favor of Supreme Court term limits, 175 stated the were in favor of universal health care, 200 were in favor of universal pre-k, and 118 were in favor of abolishing the death penalty. What portion favors universal health care, at the 95 confidence level?
Answer #1

Sample size (n) = 325

Those in favor of universal health care(x) = 175

Sample proportion = x/n = 175/325 = 0.5385

Confidence interval(in %) = 95

z @ 95% = 1.96

Since we know that

Required confidence interval = (0.5385-0.0543, 0.5385+0.0543)

Required confidence interval = (0.4842, 0.5928)

Please hit thumps up if the answer helped you.